OpenStudy (anonymous):

focus (0, 12) directrix (y=-1/2)

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

Let (x,y) be any point on the locus. If it is parabola then distance of (x,y) from focus=distance from directrix. \[\sqrt{\left( y-12 \right)^2+\left( x-0 \right)^2}=\frac{ y+\frac{ 1 }{2 } }{ \sqrt{0^2+\left( 1 \right)^2} }\] \[\left( y-12 \right)^2+x^2=\frac{ \left( y+\frac{ 1 }{ 2 } \right)^2 }{ 1 }\] simplify it.